OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

92

OpenStudy (anonymous):

267-1-2=264 267/3=88 88+1+2=92

OpenStudy (anonymous):

let the integer be x-1,x,x+1 then by the problem x-1+x+x+1=267 or 3x=267 or x=267/3 =89 hence the largest is 89+1=90

OpenStudy (anonymous):

88+1+2=91 not 92

OpenStudy (anonymous):

Sorry! 267/3=89 89+1+2=92